A bag has 4 yellow balls and 6 red balls only. One ball is chosen at random and not replaced. A second ball is then chosen at random. Find the probability that both balls are yellow.

A. 2/15

B. 4/25

C. 2/5

D. 11/15


1
Expert's answer
2021-11-19T11:07:30-0500

P(the 1st ball is yellow) "= \frac{4}{4+6} = \frac{4}{10} = \frac{2}{5}"

P(the 2nd ball is yellow) "= \frac{3}{9} = \frac{1}{3}"

The required probability "= \frac{2}{5} \times \frac{1}{3} = \frac{2}{15}"

Answer: A. 2/15